40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Mar 6 01:19:19 2020, gautam, Update, IOO, Excess laser frequency noise MCF.pdfExcessFreqNoise.pdfALSnoise.pdf
    Reply  Fri Mar 6 16:33:11 2020, gautam, Update, IOO, Excess laser frequency noise 
Message ID: 15259     Entry time: Fri Mar 6 01:19:19 2020     Reply to this: 15260
Author: gautam 
Type: Update 
Category: IOO 
Subject: Excess laser frequency noise 

Summary:

Sometime between 1PM and 6PM on Tuesday, excess laser frequency noise shows up in MCF at around 800 Hz, as shown in Attachment #1. Sigh.

Details:

While I show the MCF spectrum here, I confirmed that this noise is not injected by the IMC loop (with the PSL shutter closed, and the IMC servo board disconnected from the feedback path to the NPRO, the PMC error and control points still show the elevated noise, see Attachment #2). I don't think the problem is from the PMC loop - see Attachment #3 which is the ALS beat out-of-loop noise with the PMC unlocked (the PSL beam doesn't see the cavity before it gets to the ALS setup, and we only actuate on the cavity length for that loop, so this wasn't even really necessary).

Was there some work on the PSL table on Tuesday afternoon that can explain this

Attachment 1: MCF.pdf  122 kB  | Hide | Hide all
MCF.pdf
Attachment 2: ExcessFreqNoise.pdf  64 kB  | Hide | Hide all
ExcessFreqNoise.pdf
Attachment 3: ALSnoise.pdf  42 kB  Uploaded Fri Mar 6 10:01:06 2020  | Hide | Hide all
ALSnoise.pdf
ELOG V3.1.3-